# Revolutionizing Finance with LLMs: An Overview of Applications and Insights

### v2

@@ 4.1 Datasets
Figure 5: We conducted sentiment analysis on 970 data points from the FiQA-SA task set[128]. By using GPT-4, we achieve 79\$ accuracy rate.
@@
Figure 8: We conducted stock price prediction on 1,470 data points from the BigData task set[128]. By using GPT-4, we achieve 51\$ accuracy rate.
Figure 9: Fraud detection on PaySim simulates mobile money transactions dataset [82] using GPT-4: 5 out of 5 Correct. Green highlight indicates normal transactions; Yellow highlight indicates suspicious/fraudulent transactions.
Evaluating Sentiment in Financial News: The task of discerning sentiment in financial news is a paramount concern within the financial analytics community, as underscored in seminal works by Araci  [5] and Yang et al.   [131]. This endeavor seeks to meticulously interpret the sentiment embedded in financial narratives. Adhering to the established FLUE framework  [110], this study employs two prominent datasets: the Financial Phrase Bank (FPB) dataset  [89] and FiQA-SA  [87]. The FPB dataset is a collection of financial news excerpts, each meticulously annotated by field experts with a sentiment classification: positive, negative, or neutral. Conversely, FiQA-SA serves as an expansive dataset predominantly utilized for the sentiment quantification of English-language financial reporting and microblogging content, using a nuanced sentiment intensity scale ranging from -1 to 1, where a value of 1 epitomizes the most positive sentiment.
Identifying Named Entities in Finance: The goal of this task is to pinpoint key financial entities, including individuals, organizations, and locations. These entities are crucial for developing financial knowledge graphs. The NER dataset  [4] serves as the basis for this task, featuring sentences from financial agreements filed with the U.S. Securities and Exchange Commission and includes entities categorized as LOCATION, ORGANISATION, and PERSON.
Financial Question Answering: This task involves automatically responding to financial queries based on provided data. For this, two datasets are employed: FinQA  [20] and ConvFinQA  [21]. FinQA offers pairs of questions and answers, annotated by specialists, along with associated earnings reports from S&P 500 companies. ConvFinQA extends this by including multi-turn dialogues over these earnings reports.
Predicting Stock Movements: Recognized as a critical financial task, predicting stock movements can be invaluable in practical applications like investment strategy formulation. This task is approached as a binary classification challenge, following the methodology of prior research  [112]. It involves forecasting the direction of stock price movements based on historical prices and relevant tweets. Movements above 0.55% are considered positive, while those below -0.5% are deemed negative. For this analysis, one widely used datasets is utilized: BigData22  [112].
@@ 6 Limitation and Future work
The limitations of LLMs are evident in areas such as optimization and quantitative trading. While they can assist in identifying market sentiments, LLMs cannot directly engage in computational tasks. Their role is more auxiliary, aiding in sentiment analysis which then feeds into existing models that handle quantitative variables [123]. This indicates that LLMs, as of now, are not standalone solutions for computational finance tasks but rather powerful tools for augmenting existing models.
For future work, there is immense potential in integrating LLMs with advanced quantitative models. One promising direction could be the development of hybrid systems that combine the text processing prowess of LLMs with sophisticated quantitative trading algorithms [135, 22]. Another area could be enhancing the interpretability and reliability of LLMs outputs in financial contexts, ensuring that the insights generated are not only accurate but also actionable. Moreover, exploring the application of LLMs in predictive analytics for market trends, based on historical data and current events, can open new avenues in financial forecasting. This integration of qualitative and quantitative analysis could revolutionize how financial markets are analyzed and traded [39].

### v3

@@ 4.1 Datasets
Figure 5: We conducted sentiment analysis on 970 data points from the FiQA-SA task set[xie2023pixiu]. By using GPT-4, we achieve 79\$ accuracy rate.
@@
Figure 8: We conducted stock price prediction on 1,470 data points from the BigData task set[xie2023pixiu]. By using GPT-4, we achieve 51\$ accuracy rate.
Figure 9: Fraud detection on PaySim simulates mobile money transactions dataset [lopez2016paysim] using GPT-4: 5 out of 5 Correct. Green highlight indicates normal transactions; Yellow highlight indicates suspicious/fraudulent transactions.
Evaluating Sentiment in Financial News: The task of discerning sentiment in financial news is a paramount concern within the financial analytics community, as underscored in seminal works by Araci  [araci2019finbert] and Yang et al.   [yang2020finbert]. This endeavor seeks to meticulously interpret the sentiment embedded in financial narratives. Adhering to the established FLUE framework  [shah2022flue], this study employs two prominent datasets: the Financial Phrase Bank (FPB) dataset  [malo2014good] and FiQA-SA  [maia201818]. The FPB dataset is a collection of financial news excerpts, each meticulously annotated by field experts with a sentiment classification: positive, negative, or neutral. Conversely, FiQA-SA serves as an expansive dataset predominantly utilized for the sentiment quantification of English-language financial reporting and microblogging content, using a nuanced sentiment intensity scale ranging from -1 to 1, where a value of 1 epitomizes the most positive sentiment.
Identifying Named Entities in Finance: The goal of this task is to pinpoint key financial entities, including individuals, organizations, and locations. These entities are crucial for developing financial knowledge graphs. The NER dataset  [alvarado2015domain] serves as the basis for this task, featuring sentences from financial agreements filed with the U.S. Securities and Exchange Commission and includes entities categorized as LOCATION, ORGANISATION, and PERSON.
Financial Question Answering: This task involves automatically responding to financial queries based on provided data. For this, two datasets are employed: FinQA  [chen2021finqa] and ConvFinQA  [chen2022convfinqa]. FinQA offers pairs of questions and answers, annotated by specialists, along with associated earnings reports from S&P 500 companies. ConvFinQA extends this by including multi-turn dialogues over these earnings reports.
Predicting Stock Movements: Recognized as a critical financial task, predicting stock movements can be invaluable in practical applications like investment strategy formulation. This task is approached as a binary classification challenge, following the methodology of prior research  [soun2022accurate]. It involves forecasting the direction of stock price movements based on historical prices and relevant tweets. Movements above 0.55% are considered positive, while those below -0.5% are deemed negative. For this analysis, one widely used datasets is utilized: BigData22  [soun2022accurate].
@@ 6 Limitation and Future work
The limitations of LLMs are evident in areas such as optimization and quantitative trading. While they can assist in identifying market sentiments, LLMs cannot directly engage in computational tasks. Their role is more auxiliary, aiding in sentiment analysis which then feeds into existing models that handle quantitative variables [wang2023alphagpt]. This indicates that LLMs, as of now, are not standalone solutions for computational finance tasks but rather powerful tools for augmenting existing models.
For future work, there is immense potential in integrating LLMs with advanced quantitative models. One promising direction could be the development of hybrid systems that combine the text processing prowess of LLMs with sophisticated quantitative trading algorithms [yu2023finme, ciuriak2023trading]. Another area could be enhancing the interpretability and reliability of LLMs outputs in financial contexts, ensuring that the insights generated are not only accurate but also actionable. Moreover, exploring the application of LLMs in predictive analytics for market trends, based on historical data and current events, can open new avenues in financial forecasting. This integration of qualitative and quantitative analysis could revolutionize how financial markets are analyzed and traded [ge2023openagi].